We have a dataset with n= 10pairs of observations(xi,yi), and n∑i=1xi= 683,n∑i=1yi= 813,n∑i=1x2i= 47,405,n∑i=1xiyi= 56,089,n∑i=1y2i= 66,731.What is an approximate 99% confidence interval for the mean response atx0= 90?
Answer #1

Solution :

n= 10
Σx= 683
Σx2 = 47405
Σy = 813
Σy2 = 66731
Σxy= 56089
SSx=Σx2-(Σx)2/n= 756.1
SSy=Σy2-(Σy)2/n= 634.1
SP=Σxy-(ΣxΣy)/n= 561.1
b1= SP/Sxx = 0.7421
b=(Σy-bo*Σx)/n= 30.6147
SSE =Syy-(Sxy)2/Sxx= 217.71
σ̂2=SSE/(n-2)= 27.21363
σ̂=√σ̂2= 5.216668

from above:

n= 10
bo= 30.6147335
b1= 0.7421
sxy =√MSE= 5.2167
Sxx=(n-1)sx^2= 756.1
x̅ = 68.3
predcited value at X=90: =30.6147+90*0.7421 = 97.4035
std error confidence interval= s*√(1/n+(x0-x̅)2/Sxx) = 4.4351
for 99 % CI value of t= 3.3550
margin of error E=t*std error                            = 14.8796
lower confidence bound=sample mean-margin of error = 82.5239
Upper confidence bound=sample mean+margin of error= 112.2831

from above 99% CI =82.5239 , 112.2831

Please like if it helps me please please

Thank you so much